The problem is:
Let x and y represent two real numbers. Write an algebraic expression to denote the quotient obtained when the product of the two numbers is divided by their sum. Show your work for credit.

Answer by rothauserc(4718)

xy / (x+y) = q + r where q is the quotient and r is the remainder
therefore
xy / (x+y) = 1 +(xy -x -y)
q = 1, r = (xy -x -y)